Consider the following abbreviated financial statements for Pinghua:

PINGHUA
2014 and 2015 Partial Balance Sheets

Assets

Liabilities and Owners’ Equity

2014

2015

2014

2015

  Current assets

$

1,850

$

2,100

Current liabilities

$

740

$

860

  Net fixed assets

7,850

9,120

Long-term debt

4,000

4,300

Equity

4,960

6,060

PINGHUA
2015 Income Statement Info

  Sales

$

11,300

  Costs

5,500

  Depreciation

1,000

  Interest paid

200

The tax rate is 35%. Long term debt trades at 112% of par. The firm has 500 shares outstanding. Free Cash Flow to the Firm and the Free Cash Flow to Equity are both expected to grow at 2% forever. The weighted average cost of capital is 16%. The cost of equity is 24%.

Question 9

What is the Net Income for 2015? $______

Hint: Construct the Income Stmt using the partial info. You are almost there, you just have to compute the taxable income and subtract the income tax.

What is the Free Cash Flow to the Firm for 2015? $______

Hint: FCFF = ? = EBIT-Tax+D-chgNWC-Capex where Capex=chgNFA+D

What is the Value of the Firm (Assets)? $______

Hint: FCFF / (WACC -g)

What is the stock price per share using the discounted FCFF valuation? $______

Hint: Take the previous answer, subtract the market value of Debt, that will produce the value of the Equity. Then divide by the number of shares.

What is the Free Cash Flow to Equity for 2015? $______

Hint: The easiest way is FCFE = FCFF - interest + chgDebt or follow the original FCFE equation.

What is the stock price per share using the discounted FCFE valuation? $______

Hint: FCFE / (RE - g) gives you the value of Equity, then divide by the number of shares.
Answer #1
Q.9.
PINGHUA
2015 Income Statement Amt $
Salces $          11,300
Costs $            5,500
Dereciation $            1,000
EBIT $            4,800
Interest paid $                200
EBT $            4,600
Tax @35% $            1,610
Net Income $            2,990
ans a. So Net Income for 2015=$2,990
FCFF Calculation
Balance sheet partial data 2014 2015
1 Current Asset $            1,850 $            2,100
2 Less Current Liabilities $                740 $               860
Working Capital =1-2= $            1,110 $            1,240
Increase in Working Capital $               130
3 Net Fixed Asset $            7,850 $            9,120
4 Increase in NFA $            1,270
5 Add Depreciation                 1,000
Additional Capital Spending =4+5= $            2,270
FCFF =EBIT-Tax +Depreciation-Capital expenditure-Change in NWC
so FCFF=4800-1610+1000-2270-130 = $       1,790.00
Ans b. So Free Cash flow to Firm=$1,790
The perpetual growth rate is g=2%
WACC=16%
Ans c. So Value of Firm FCFF/(WACC-g)=1790/(16%-2%)= $    12,785.71
Book value of debt in 2015 =4300
Market value of debt @112% of par in 2015= $            4,816
Now value of firm=$12785.71
Value of Equity =Value of firm-Market Value of debt= $       7,969.71
No of shares outstanding = 500
Ans d. Value per share =7969.71/500= $            15.94
Balance sheet partial data 2014 2015
Long term debt $            4,000 $            4,300
Additional debt raised $               300
FCFE calculation
FCFE=FCFF-Interest +change in Debt
FCFE=1790-200+300=1890
Ans e So Free cash flow to Equity =$1,890
Cost of equity =k=24%
The perpetual growth rate is g=2%
Value of Equity =FCFE/(k-g)=1890/(24%-2%)= $       8,590.91
No of shares outstanding = 500
Ans f Stock price/share=8590.91/500= $            17.18